n December (month 24) of year 2 of operations, Big Al’s produced and sold 32,675 pizzas, consisting of 30,570 meat pizzas and 2,105 veggie pizzas. The budgeted sales price for meat pizzas was $5.00, and for veggie pizzas, $5.25. The estimated production and sales during December was 31,678 meat pizzas and 2,595 veggie pizzas.

Required:
A. Compute the price and volume variances for sales, assuming that Big Al’s sold all pizzas produced for $137,565 (meat) and $9,051 (veggie). What might explain these variances?
B. Compute the price and quantity variances for direct materials for each type of pizza, assuming that Big Al’s paid $29,093 for 32,325 units of raw material for meat pizzas and $1,453 for 2,401 units of raw material for veggie pizzas.(A unit consists of dough shell, sauce, cheese, meat or veggies, and assembly materials.) In addition, 30,995 units were used to produce meat pizzas, and 2,149 units were used to produce veggie pizzas. How would these variances be interpreted? What might explain these variances? Would you consider them to be large enough to be important?
C. Compute the labor rate and efficiency variances, assuming that Big Al’s paid $71,350 in labor costs for 7,150 hours of labor for meat pizzas and $6,425 in labor costs for 650 hours of labor for veggie pizzas. How would these variances be interpreted? What might explain them? Would you consider them large enough to be important?
Answer #1

A. Material price Variance= (SP-AP)*AQ

Veggie Pizza= (5.25-3.49)*2595 =4567.2 Favorable

Meat Pizza= (5-4.34)31678 =20907.48 Favorable

Volume Variance= (AQ-SQ)*SP

Veggie Pizza= (2595-2105)*5.25 =2572.5 Favorable

Meat Pizza= (31678-30570)5.0 =5540 Favorable

B.